# Jet-like structures in low-mass binary neutron star merger remnants

### v1

@@ III.2 SMNS remnants
The BZ mechanism cannot operate without an ergosphere Ruiz et al. (2020b), however the rotation of the magnetized stars twists the frozen-in magnetic field lines in much the same fashion, forming same kind of tightly wound helical magnetic structure seen in the IR2.70 case (see Fig. 2 lower right). The Poynting flux generated by the rotating coiled magnetic field accelerates an outflow wind of gas, again forming an evacuated funnel, although unlike the BH case, there remains a high density region of $\rho_{0}\sim 10^{11}\textup{g}\,\textup{cm}^{-3}$ at the base of the funnel extending to $\sim 20$km above the NS poles (see Fig. 4 bottom panel, Fig. 5 top row and Fig. 10 top panel). The funnel itself is also narrower (half-opening angle of $\sim 25^{\circ}$) and much more polluted with baryonic gas than for the IR2.70 BH case (Fig. 5 top row).
@@ III.4 Effects of initial NS spin: Comparison between spinning and irrotational cases
The first obvious impact of initial NS spin is that it delays the inspiral of the stars by $\sim 2$ orbits (Fig. 6): the well-known “hang-up” effect due to spin-orbit coupling Campanelli et al. (2006); Tsokaros et al. (2019b). The additional angular momentum in the system due to the spin of the initial NSs results in larger amounts of dynamical ejecta for the same initial mass and a larger and more diffuse debris disk (see Table 3). The EM luminosity prior to merger is substantially larger for the spinning cases than for the irrotational cases, as the frozen-in magnetic field is subject to more winding. Following merger, the EM luminosities remain consistently larger for spinning compared to irrotational cases, although only by a factor of $\sim 1.25-2$. We also see that the post-merger higher frequency component of the GW signal is supressed in the spinning cases compared to the irrotational cases (see Sec. III.7) although for a merger at a distance $\sim 50$Mpc this is unlikely to be detectable.
@@ III.6 Magnetic energy amplification
The total magnetic energy, including the EM outflow from the simulation box, is shown in Fig. 12. Immediately after merger the magnetic energy increases by over an order of magnitude within a few ms, consistent with amplification via the KHI instability at the shear surface Kiuchi et al. (2015) (with KH vortices visible). The relative amplification in the first 1.5ms is shown in Fig. 13 (left panel) with the black lines indicating fits to the initial linear regime for the irrotational (IR) and spinning (SP) cases respectively. The SP cases exhibit slower amplification ($\gamma_{\textup{KH}}\sim 1.09$) compared to the irrotational cases ($\gamma_{\textup{KH}}\sim 1.38$), which we can attribute to the reduced speed difference across the shear surface (shown in the right panel of Fig. 13). Note however that the magnetic energy growth rate $2\gamma_{\textup{KH}}\sim 2\textup{ms}^{-1}$ is several orders of magnitude lower than the expected instability growth rate from linear perturbation theory of $\sigma_{\textup{KH}}\sim 10^{2}\textup{ms}^{-1}$ from $\Delta v\sim 0.1c,d\sim 400\rm m$. After the KHI growth terminates at $\sim 2-6$ms, with a maximum magnetic field strength of $\sim 10^{17}$, the magnetic energy falls. This decline (also seen in higher resolution simulations Palenzuela et al. (2022)) may be due to the conversion of magnetic to kinetic energy in accelerating the ejecta or to the collapse of an unstable magnetic configuration. However, further study is needed to precisely determine the dynamics.
@@ III.7 Gravitational waves
Figure 15: Characteristic frequencies of the post-merger GW signal for the different NS remnants vs initial total binary gravitational mass. Error bars denote uncertainties taken from the width of the fitted Gaussian peaks. The solid orange and dashed pink lines denote the $f_{0}$ quasi-radial frequency, as measured by fitting a Lorentzian peak to the Fourier spectrum of the oscillations of the minimum lapse $\alpha_{\textup{min}}$, for the irrotational cases and spinning cases respectively. Also plotted are empirical quasi-universal relations given in several previous GRHD (no magnetic fields) numerical studies.
@@ IV Conclusions
For the SMNS cases we also observe the formation of a partially evacuated funnel (half-opening angle $\sim 25^{\circ}$) above the poles of the NS remnants. We observe a helical magnetic field within this funnel region that is actually stronger than in the HMNS/BH case, a mildly relativistic outflow, and find that the SMNS cases produce a high fluid and EM luminosity ($\sim 10^{53}\textup{erg}\;\textup{s}^{-1}$) for a longer duration than the HMNS case. From the ejecta mass ($M_{\textup{esc}}\gtrsim 10^{-2}M_{\odot}$) and average ejecta velocity ($\langle v_{\textup{eje}}\rangle\sim 0.2c$) we estimate the bolometric luminosity for the associated kilonova as $L_{\textup{knova}}\sim 10^{41}\textup{erg}\;\textup{s}^{-1}$, and both the EM and kilonova luminosity are broadly consistent with observations of sGRBs Li et al. (2016); Beniamini et al. (2020) and kilonovae Ascenzi et al. (2019); Rastinejad et al. (2021); Villar et al. (2017), although towards the high end of the observable range. However, the gas density in the funnel is also significantly larger and this baryon pollution means the outflow is less magnetically dominated (i.e. $b^{2}/(2\rho_{0})$ is smaller compared to the HMNS case) limiting the magnetic energy per unit mass available for accelerating the gas to ultrarelativistic speeds.
In addition, we note that for outflows from magnetized NS remnants we do not have the same kind of evidence for efficient acceleration and ultrarelativistic jet formation that we have for BZ-powered jets from BHs with magnetized accretion disks. Our results suggest one cannot simply assume that magnetic energy at the base of the outflow is efficiently converted to kinetic energy asymptotically, as for idealized semi-analytic models of ideal MHD jets, and that in real dynamical system with a large amount of baryon loading significant amounts of energy may be lost via mixing between the low density funnel and the higher density debris torus. However, we emphasise that further simulations on larger spatial scales, akin to e.g. Gottlieb et al. (2022a), are needed to fully resolve this question. Therefore while we note that we observe jet-like structures in the outflow from the SMNS cases which meet our essential criteria for an incipient jet, we cannot conclude they will actually produce a true ultrarelativistic jet as required for the observed $\gamma-$ray emission or correspond to the central engines for sGRBs.
@@ Acknowledgements
We thank members of our Illinois Relativity Undergraduate Research Team (Nawaf Aldrees, Jonah Doppelt, Rohan Narasimhan, Yuan Liang, Eric Yu) for assistance with some of the 3D visualizations. This work was supported in part by National Science Foundation (NSF) Grants No. PHY-2308242, No. OAC-2310548 and No. PHY-2006066 to the University of Illinois at Urbana-Champaign. M.R. acknowledges support by the Generalitat Valenciana Grant CIDEGENT/2021/046 and by the Spanish Agencia Estatal de Investigación (Grant PID2021-125485NB-C21). A.T. acknowledges support from the National Center for Supercomputing Applications (NCSA) at the University of Illinois at Urbana-Champaign through the NCSA Fellows program. This work used Stampede2 at TACC and Anvil at Purdue University through allocation MCA99S008, from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research also used Frontera at TACC through allocation AST20025. Frontera is made possible by NSF award OAC-1818253.

### v2

@@ III.2 SMNS remnants
The BZ mechanism cannot operate without an ergosphere Ruiz et al. (2020b), however the rotation of the magnetized stars twists the frozen-in magnetic field lines in much the same fashion, forming same kind of tightly wound helical magnetic structure seen in the IR2.70 case (see Fig. 2 lower right). The Poynting flux generated by the rotating coiled magnetic field accelerates an outflow wind of gas, again forming an evacuated funnel, although unlike the BH case, there remains a high density region of $\rho_{0}\sim 10^{11}\textup{g}\,\textup{cm}^{-3}$ at the base of the funnel extending to $\sim 20$km above the NS poles (see Fig. 4 bottom panel, Fig. 6 top row and Fig. 11 top panel). The funnel itself is also narrower (half-opening angle of $\sim 25^{\circ}$ at the base) and much more polluted with baryonic gas than for the IR2.70 BH case (Fig. 6 top row).
@@ III.4 Effects of initial NS spin: Comparison between spinning and irrotational cases
The first obvious impact of initial NS spin is that it delays the inspiral of the stars by $\sim 2$ orbits (Fig. 5): the well-known “hang-up” effect due to spin-orbit coupling Campanelli et al. (2006); Tsokaros et al. (2019b). The additional angular momentum in the system due to the spin of the initial NSs results in larger amounts of dynamical ejecta for the same initial mass and a larger and more diffuse debris disk (see Table 3). The EM luminosity prior to merger is substantially larger for the spinning cases than for the irrotational cases, as the frozen-in magnetic field is subject to more winding. Following merger, the EM luminosities remain consistently larger for spinning compared to irrotational cases, although only by a factor of $\sim 1.25-2$. Our results also show that the half-opening angle of the Poynting luminosity per solid angle at large distances is consistently slightly larger for the spinning vs. the irrotational cases (see Fig. 7 center and right panels for the representative $M=2.40M_{\odot}$ cases), giving a tantalizing hint that the additional angular momentum from spinning progenitors may alter the outflow morphology. We hope to explore this further in a future work. Finally, we see that the post-merger higher frequency component of the GW signal is supressed in the spinning cases compared to the irrotational cases (see Sec. III.7).
@@ III.6 Magnetic energy amplification
The total magnetic energy, including the EM outflow from the simulation box, is shown in Fig. 13. Immediately after merger the magnetic energy increases by over an order of magnitude within a few ms, consistent with amplification via the KHI instability at the shear surface Kiuchi et al. (2015) (with KH vortices visible). The relative amplification in the first 1.5ms is shown in Fig. 14 (left panel) with the black lines indicating fits to the initial linear regime for the irrotational (IR) and spinning (SP) cases respectively. The SP cases exhibit slower amplification ($\gamma_{\textup{KH}}\sim 1.09$) compared to the irrotational cases ($\gamma_{\textup{KH}}\sim 1.38$), which we can attribute to the reduced speed difference across the shear surface (shown in the right panel of Fig. 14). Note however that the magnetic energy growth rate $2\gamma_{\textup{KH}}\sim 2\textup{ms}^{-1}$ is several orders of magnitude lower than the expected instability growth rate from linear perturbation theory of $\sigma_{\textup{KH}}\sim 10^{2}\textup{ms}^{-1}$ from $\Delta v\sim 0.1c,d\sim 400\rm m$. After the KHI growth terminates at $\sim 2-6$ms, with a maximum magnetic field strength of $\sim 10^{17}$, the magnetic energy falls. This decline (also seen in higher resolution simulations Palenzuela et al. (2022); Aguilera-Miret et al. (2023)) may be due to the conversion of magnetic to kinetic energy in accelerating the ejecta or to the collapse of an unstable magnetic configuration. However, further study is needed to precisely determine the dynamics.
@@ III.7 Gravitational waves
Figure 16: Characteristic frequencies of the post-merger GW signal for the different NS remnants vs initial total binary gravitational mass. Error bars denote uncertainties taken from the width of the fitted Gaussian peaks. The solid orange and dashed pink lines denote the $f_{0}$ quasi-radial frequency, as measured by fitting a Lorentzian peak to the Fourier spectrum of the oscillations of the minimum lapse $\alpha_{\textup{min}}$, for the irrotational cases and spinning cases respectively. Also plotted are empirical quasi-universal relations given in several previous GRHD (general relativistic hydrodynamics) numerical studies without magnetic fields.
@@ IV Conclusions
For the SMNS cases we also observe the formation of a partially evacuated funnel (half-opening angle $\sim 25^{\circ}$) above the poles of the NS remnants. We observe a helical magnetic field within this funnel region that is actually stronger than in the HMNS/BH case, a mildly relativistic outflow, and find that the SMNS cases produce a high fluid and EM luminosity ($\sim 10^{53}\textup{erg}\;\textup{s}^{-1}$) for a longer duration than the HMNS case. From the ejecta mass ($M_{\textup{esc}}\gtrsim 10^{-2}M_{\odot}$) and average ejecta velocity ($\langle v_{\textup{eje}}\rangle\sim 0.2c$) we estimate the bolometric luminosity for the associated kilonova as $L_{\textup{knova}}\sim 10^{41}\textup{erg}\;\textup{s}^{-1}$, broadly consistent with observations Ascenzi et al. (2019); Rastinejad et al. (2021); Villar et al. (2017). The EM luminosity may be consistent with the higher end of observed $\gamma-$ray luminosities from sGRBs Li et al. (2016); Beniamini et al. (2020), or in excess, depending on uncertainties about the $\gamma-$ray prompt emission mechanism and the efficiency of generating $\gamma-$ray photons.
The gas density in the funnel is also significantly larger compared to the BH case, and this baryon pollution means the outflow is less magnetically dominated (i.e. $b^{2}/(2\rho_{0})$ is smaller) limiting the magnetic energy per unit mass available for accelerating the gas to ultrarelativistic speeds.
In addition, we note that for outflows from magnetized NS remnants we do not have the same kind of evidence for efficient acceleration and ultrarelativistic jet formation that we have for BZ-powered jets from BHs with magnetized accretion disks. Our results suggest one cannot simply assume that magnetic energy at the base of the outflow is efficiently converted to kinetic energy asymptotically, as for idealized semi-analytic models of ideal MHD jets, and that in real dynamical system with a large amount of baryon loading significant amounts of energy may be lost via mixing between the low density funnel and the higher density debris torus (which may also reduce the $\gamma-$ray conversion efficiency). However, we emphasise that further simulations on larger spatial scales, akin to e.g. Gottlieb et al. (2022a), are needed to fully resolve this question. Therefore while we note that we observe jet-like structures in the outflow from the SMNS cases which meet our essential criteria for an incipient jet, we cannot conclude they will actually produce a true ultrarelativistic jet as required for the observed $\gamma-$ray emission or correspond to the central engines for sGRBs.
@@
Movies and additional 3D visualizations highlighting our simulations can be found at web (2024).
@@ Acknowledgements
We thank members of our Illinois Relativity Undergraduate Research Team (Nawaf Aldrees, Jonah Doppelt, Rohan Narasimhan, Yinuan Liang, Eric Yu) for assistance with some of the 3D visualizations. This work was supported in part by National Science Foundation (NSF) Grants No. PHY-2308242, No. OAC-2310548 and No. PHY-2006066 to the University of Illinois at Urbana-Champaign. M.R. acknowledges support by the Generalitat Valenciana Grant CIDEGENT/2021/046 and by the Spanish Agencia Estatal de Investigación (Grant PID2021-125485NB-C21). A.T. acknowledges support from the National Center for Supercomputing Applications (NCSA) at the University of Illinois at Urbana-Champaign through the NCSA Fellows program. This work used Stampede2 at TACC and Anvil at Purdue University through allocation MCA99S008, from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research also used Frontera at TACC through allocation AST20025. Frontera is made possible by NSF award OAC-1818253.